# arXiv:2603.01030v1[math.NA] 01 Mar 2026

### 01 Mar 2026 v1

@@ 5.2 Improved consistency error estimate
$\|h_{\mathcal{T}}^{-1}\left(r\bm{v}-\Pi_{\mathrm{RT}0}^{\mathrm{axi}}(r\bm{v})\right)\|_{L^{2}_{-1}(\Omega)}\lesssim\|\bm{v}\|_{\bm{V}}.$
@@
For triangles that are not of type 1 or type 2, the estimate follows from the usual $\mathrm{RT}_{0}$ interpolation estimates (scaled by $r_{\text{max}}/r_{\text{min}}\approx 1$) with order $h$ instead of $h^{3/2}$. Hence, the proof concentrates on type 1 or type 2 triangles, such that the modified basis functions are involved.
@@
$\|\mathbb{P}(\Delta_{\text{axi}}\bm{u})\circ(\mathbb{I}-\Pi)\|_{\bm{V}^{\star}}\lesssim\|h_{\mathcal{T}}\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}_{1}}.$
@@
$(\mathbb{P}(\Delta_{\text{axi}}\bm{u}),r\bm{w}_{h}-\Pi(r\bm{w}_{h}))_{L^{2}}$ $\leq\|h_{\mathcal{T}}\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}_{1}}\|h_{\mathcal{T}}^{-1}(1-\Pi)(r\bm{w}_{h})\|_{L^{2}_{-1}}$
$\lesssim\|h_{\mathcal{T}}\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}_{1}}\|\bm{w}_{h}\|_{\bm{V}}.$
@@ 6.2 Example 2: smooth example
We consider an example with smooth data, where we expect all methods to show optimal convergence rates. Specifically, we choose
@@ 6.3 Example 3: data in $L^{2}_{1}(\Omega)\setminus L^{2}(\Omega)$
However, it is interesting that the critical value for the modified reconstruction operators is much smaller than that for the classical reconstruction operators for the same quadrature rule and that seems to be related to the properties of the reconstruction operator. More importantly, it can be observed that the classical reconstruction operator $\Pi_{\mathrm{RT}0}$ yields much larger errors than the Galerkin method for moderate $\nu$.

### 13 Mar 2026 v2

@@ 5.2 Improved consistency error estimate
$\|h^{-1}\left(r\bm{v}-\Pi_{\mathrm{RT}0}^{\mathrm{axi}}(r\bm{v})\right)\|_{L^{2}_{-1}(\Omega)}\lesssim\|\bm{v}\|_{\bm{V}}.$
@@
For triangles that are not of type 1 or type 2, the estimate follows from the usual $\mathrm{RT}_{0}$ interpolation estimates (scaled by $r_{\text{max}}/r_{\text{min}}\approx 1$, where $r_{\text{min}}$ and $r_{\text{max}}$ denote the minimum and maximum $r$-coordinates on $T$) with order $h$ instead of $h^{3/2}$. Hence, the proof concentrates on type 1 or type 2 triangles, such that the modified basis functions are involved.
@@
$\|\mathbb{P}(\Delta_{\text{axi}}\bm{u})\circ(\mathbb{I}-\Pi)\|_{\bm{V}^{\star}}\lesssim\|h\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}_{1}}.$
@@
$(\mathbb{P}(\Delta_{\text{axi}}\bm{u}),r\bm{w}_{h}-\Pi(r\bm{w}_{h}))_{L^{2}}$ $\leq\|h\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}_{1}}\|h^{-1}(1-\Pi)(r\bm{w}_{h})\|_{L^{2}_{-1}}$
$\lesssim\|h\mathbb{P}(\Delta_{\text{axi}}\bm{u})\|_{L^{2}_{1}}\|\bm{w}_{h}\|_{\bm{V}}.$
@@ 6.2 Example 2: smooth example
This example considers smooth data, where we expect that all methods to show optimal convergence rates, namely
@@ 6.3 Example 3: data in $L^{2}_{1}(\Omega)\setminus L^{2}(\Omega)$
However, it is interesting that the critical value for the modified reconstruction operators is much smaller than that for the classical reconstruction operators for the same quadrature rule. Another interesting observation is that the error for the unmodified reconstruction operators increases when the quadrature order is increased and quadrature points move closer to the rotation axis. This behavior seems to be related to the missing vanishing-on-axis-property of the unmodified reconstruction operators and the missing $L^{2}_{1}$-regularity of the data. The modified method and the classical method do not show an increase of the error when the quadrature order is increased. More importantly, it can be observed that the classical reconstruction operator $\Pi_{\mathrm{RT}0}$ yields much larger errors than the Galerkin method for moderate $\nu$.